A dollar coaster travels 41.1 m at an angle of 40.0° above the horizontal. How far does it move horizontally and vertically?
1
Expert's answer
2019-10-25T11:02:24-0400

The horizontal distance:


"x=s\cos{40}=41.1\cos{40}=31.5\ m"

The vertical distance:


"y=s\sin{40}=41.1\sin{40}=26.4\ m"